Use the definition of limit to prove that both of the sequences { 1/√n } and { (−1)^n/√ n } converges to 0.


1
Expert's answer
2021-01-26T15:10:21-0500

To show that "\lim_{n \to \infty} \frac{1}{\sqrt n} = 0"

We show that given any "\epsilon > 0" there exist "M \in \mathbb{N}" such that "n>M \implies" "\mid \frac{1}{\sqrt n} - 0 \mid < \epsilon"

Set "M> \frac{1}{\epsilon^{2}} \implies \epsilon > \frac{1}{\sqrt M}"

So, whenever "n> M"

Consider

"\mid \frac{1}{\sqrt n} - 0 \mid = \mid \frac {1}{\sqrt n} \mid = \frac {1}{\sqrt n} < \frac {1}{\sqrt M} < \epsilon"

as desired.


To show that "\lim_{n \to \infty} \frac{(-1)^{n}}{\sqrt n} = 0"

We show that given any "\epsilon > 0" there exist "M \in \mathbb{N}" such that "n>M \implies \mid \frac{(-1)^n}{\sqrt n} - 0 \mid < \epsilon"

Set "M> \frac{1}{\epsilon^{2}} \implies \epsilon > \frac{1}{\sqrt M}"

So, whenever "n> M"

Consider

"\mid \frac{(-1)^n}{\sqrt n} - 0 \mid = \mid \frac {(-1)^n}{\sqrt n} \mid \leq \mid \frac{1}{\sqrt n}\mid = \frac {1}{\sqrt n} < \frac {1}{\sqrt M} < \epsilon"

as desired.